OpenStudy (mizuki):

Teesha is in the French club. There are 26 students in the club. The French teacher will pick 3 students at random to guide visiting students from France. What is the probability that Teesha will not be picked as a guide? 3/29 3/26 29/26 23/26

OpenStudy (mizuki):

B?

OpenStudy (mizuki):

am i correct??

OpenStudy (kropot72):

The required probability is found from: \[\large P(Teesha\ excluded)=\frac{25C3}{26C3}\]

OpenStudy (kropot72):

@Mizuki Your choice is not correct.

OpenStudy (mizuki):

how come?

OpenStudy (kropot72):

\[\large P(Teesha\ excluded)=\frac{25!}{3!22!}\times\frac{3!23!}{26!}\]